a) A rod with a length of 35 cm has a charge of -2.5x10-9C. Find the electric field magnitude at a distance of 2 cm from the center of the rod.

b) Find the electric field E at 0.02m from the surface of a sphere with radius 0.1m if it has a superficial density charge +8x10-9C/m2

1
Expert's answer
2021-10-06T08:15:42-0400

a) The magnitude of the field is


"E=\frac\lambda{2\pi\epsilon_0r}=2250\text{ V/m}."

b) The magnitude of the field is


"E=\frac{4\pi R^2\sigma}{4\pi\epsilon_0(R+b)^2}=\frac{R^2\sigma}{\epsilon_0(R+b)^2}=627\text{ V/m}."